# GEDAN: Learning the Edit Costs for Graph Edit Distance

### v2

@@ 1 Introduction and Related Work
In the last decades, several methods have been proposed and researched to assess the dissimilarity or similarity between graphs, such as diverse graph kernels or graph matching procedures (Conte et al., 2004; Vento, 2013). Among these, Graph Edit Distance (GED) (Bunke & Allermann, 1983) can be interpreted as a standard approach. Basically, GED measures the distance between two graphs in terms of the minimum cost operations required to transform one graph into the other. The considered operations typically include insertions, deletions, and substitutions of both nodes and edges. One of the most important advantages of GED is that it can be applied to virtually all types of graphs and thus, GED can be flexibly adapted to any graph-based application. However, exact computation of GED is an NP-hard problem (Bunke, 1997), which makes its application to larger graphs, or in scenarios requiring a large number of comparisons, impractical.
To address this limitation and to reduce the computational complexity of GED, several strategies for relaxing the problem have been proposed. These techniques range from binary linear programming (Justice & III, 2006), over fast suboptimal algorithms preserving classification accuracy (Neuhaus et al., 2006b; Riesen & Bunke, 2009b), to quadratic assignment models (Bougleux et al., 2017) and local search algorithms (Boria et al., 2020). More recently, diverse approaches to learn and approximate GED using Graph Neural Networks (GNNs) have been proposed (Bai et al., 2019a; Ranjan et al., 2022; Zhuo & Tan, 2022). A very recent approach (Jain et al., 2024), also based on GNNs, differs from the others in explicitly formulating the GED as a matrix optimization problem. That is, this method approximates the GED by learning the permutation that minimizes the assignment cost between the substructures of two graphs.
@@
Consequently, adapting GED to reflect meaningful, task-specific edit costs is non-trivial, since most available methods are supervised and rely on pre-computed distances. To make GED reflect functional similarity, it is necessary to learn the underlying edit costs – quantities that are not known a priori but constitute the very target of learning. While unsupervised methods such as GraphCL (You et al., 2020) or UGraphEmb (Bai et al., 2019b) can learn structural representations through augmentation, they cannot explicitly capture the actual edit costs. Recent methods such as EPIC (Heo et al., 2024) and GMSM (Pellizzoni et al., 2024) attempt to learn edit costs. EPIC employs them primarily for data augmentation, while GMSM restricts node-specific costs to the final matching step, thereby missing multi-scale granularity.
@@ 2.2 Learning the Matching
In our model the matching between two graphs $G$ and $G^{\prime}$, containing $n$ and $m$ nodes respectively, is learned by constructing a series of matrices of size $n+m]\times[m+n]$, as illustrated in Figure 1 (a).
@@
The sum $\sum_{k=0}^{K}\mathbf{M}_{k}+\mathbf{D}+\mathbf{B}$ aligns with the LSAP-based reformulation of GED (Riesen & Bunke, 2009b). In our formulation, however, we do not constrain insertions and deletions to diagonal entries; instead, a masking mechanism hides off-diagonal elements from the network to prevent them from affecting learning (see $\mathbf{Mask}$ in Figure 1 (a)).
@@ 2.4 The Training of GEDAN
In the unsupervised setting, optimization follows Eq. 7. If the weights of the functions $f_{\theta}^{(k)}$ are frozen and treated as fixed costs, GEDAN can be directly used to approximate GED, avoiding updates to the learned costs and preventing potential shortcut solutions. For consistency, the insertion and deletion costs ($a^{\ominus},a^{\oplus},b^{\ominus},b^{\oplus}$) are also kept fixed.
In the supervised setting, when target labels are available, the model leverages the edit costs learned by the functions $f_{\theta}^{(k)}$ to align GED with the target labels. This gives GEDAN great flexibility. On the one hand, it can be used to minimize GED in a standard fashion using an appropriate loss (we use Smooth L1 loss (Girshick, 2015)). On the other hand, its most interesting application arises when topological dissimilarity does not coincide with functional dissimilarity. In this case, given the target labels, edit costs are optimized through a multi-task learning strategy (Song et al., 2016) that employs multiple loss functions. Multi-loss training in metric learning has been extensively studied (Wang et al., 2019a; Schroff et al., 2015) and is known to improve generalization and the robustness of learned representations. In our model, two complementary loss functions are employed.
@@
The second loss function is introduced to supervise the downstream task (Sobal et al., 2024), which may involve either regression or classification. This loss is computed via a neural network module that processes the predicted GEDs between the query graph and a set of key graphs, and uses this information to predict the label of the query graph.
@@ 3 Experimental Evaluation
For the first experiment, we use three datasets from the TUDataset repository (Morris et al., 2020) that contain molecule graphs. The chosen datasets consist of intentionally small graphs, motivated by the need to ensure the feasibility of the supervised training required by most models. For the second and third evaluations, we employ two benchmark datasets (Wu et al., 2017; Dwivedi et al., 2023b), namely FreeSolv (Mobley & Guthrie, 2014) and BBBP (Martins et al., 2012), which describe hydration free energy and blood–brain barrier permeability, respectively.
@@ 3.2 Optimization of Edit Costs
We use the predicted GED to embed the graphs in a vector space using dissimilarity-based embeddings (Riesen & Bunke, 2009a). The prototype selection (actually required for graph embedding) is performed uniformly to ensure a fair comparison. Moreover, to ensure the robustness of our results and to avoid possible bias due to an unrepresentative selection of prototypes, we repeat the process three times (for each fold during cross-validation).
@@ 3.3 Analysis of Edit Costs
In Figure 3 (c), we examine a molecule containing functional groups associated with enhanced water solubility. In particular, the sulfonyl (–SO2–) and nitro (–NO2) groups, which have been reported to influence solubility (Goldberg & Parker, 1985; Morrison & Boyd, 1992), are correctly highlighted by all models, albeit with different levels of resolution and local precision.
@@ 4 Conclusion
In this work, we propose a novel graph-based architecture that learns edit costs end-to-end, directly aligning GED with downstream tasks and overcoming the limitations of fixed-cost approaches. Although unsupervised GED approximation can improve graph matching in principle, achieving meaningful alignment with task-specific functional dissimilarity requires learning the edit costs. Our method enables fine-grained cost optimization at the node and subgraph level, producing embeddings that capture task-relevant differences and highlighting functionally important regions within the graphs. This fine-grained alignment not only improves predictive performance but also supports interpretable analysis by revealing which graph regions contribute most to the task. Future work will focus on scaling the approach to larger graphs.
@@
Acknowledgments
This study is funded by the Swiss National Science Foundation (SNSF) - Project Nr 200020_219388.
@@ Appendix C Model Details
Training GEDAN requires substantial computational resources. We employ a NVIDIA A100 GPUs for GED approximation, while experiments on learning edit costs are conducted on a cluster with two NVIDIA H100 GPUs to optimize training time and cost. The implementation is based on PyTorch (v2.2.0) together with the PyTorch Geometric library (v2.4.0), with hardware acceleration provided by CUDA (v12.1). The source code is publicly available at https://github.com/MonsieurSolver/GEDAN.
@@ C.4 Gumbel-Sinkhorn
To obtain a good $\mathbf{P}$, we modify the version of the original Gumbel-Sinkhorn network (Mena et al., 2018; Shah et al., 2024), using the Truncated Gumbel (Neamah & Qasim, 2021) to reduce the gradient variance and add stability. We then sample with the Gumbel-Max Trick (Huijben et al., 2023) while maintaining differentiability. Finally we stabilize the normalization of Sinkhorn (Sinkhorn, 1964) with an early stop, improving efficiency and limiting gradient vanishing (LeCun et al., 1998).
@@ C.5 Influence of Depth and Inference Dynamics
This improvement is achieved in an unsupervised way, without direct supervision on the GED values, showing that $\text{GEDAN}(\lambda=1)$ can internally structure representations that aligns with true graph distances. Further examples are shown in Figures 121314 and 15.
@@ D.1 The Setting Up of the Configurations
    The first configuration serves as a baseline, where all operations—node and edge insertions, deletions, and substitutions—have uniform costs set to 1, as commonly assumed in several models.
@@
For completeness, we report the statistics of GED values for each configuration and dataset considered. Recent work indicates that GED benchmarks may have some bias related to isomorphism, or that this bias may be easily introduced during the training phase. In our case, since the main goal is to compare our model with those in the literature, we acknowledge this potential criticality but leave it explicitly as a cue for future work. However, it should be emphasized that all datasets used are public and thus available for further analysis.
@@ D.2 Results For Each Configuration
A comparative analysis across different configurations confirms the absence of a universally superior model, in agreement with the no-free-lunch theorem (Wolpert & Macready, 1997). The best performance is highlighted in boldface, while the second-best is underlined. For EUGENE (Bommakanti et al., 2024), we report the standard derivation, as we applied folding to accelerate the approximation using fewer graphs. In contrast, for the approximation proposed by Fischer et al. (2017), we approximated the entire dataset directly, since this method is efficient and, like EUGENE, does not require any training.
@@ Appendix E Details of Edit Cost Optimization
FreeSolv (Mobley & Guthrie, 2014) BBBP (Martins et al., 2012) N. graphs Avg. nodes Avg. edges N. graphs Avg. nodes Avg. edges 642 8.72 16.78 1734 20.84 44.88
FreeSolv (Mobley & Guthrie, 2014) is a small dataset that collects hydration free energies (HFE), both experimental and calculated, for small molecules in water. It is commonly used for the prediction of molecular properties in the field of computational chemistry. The choice of this dataset is due to its small size, especially in application contexts one does not have a lot of data, so we verified that the model would work in this context.
@@
E.1 Loss and Training Strategy
@@
E.2 Edit Cost Optimization Effect

### v3

@@ 1 Introduction and Related Work
In the last decades, several methods have been proposed and researched to assess the dissimilarity or similarity between graphs, such as diverse graph kernels or graph matching procedures (Conte et al., 2004; Vento, 2013). Among these, Graph Edit Distance (GED) (Bunke and Allermann, 1983) can be interpreted as a standard approach. Basically, GED measures the distance between two graphs in terms of the minimum cost operations required to transform one graph into the other. The considered operations typically include insertions, deletions, and substitutions of both nodes and edges. One of the most important advantages of GED is that it can be applied to virtually all types of graphs and thus, GED can be flexibly adapted to any graph-based application. However, exact computation of GED is an NP-hard problem (Bunke, 1997), which makes its application to larger graphs, or in scenarios requiring a large number of comparisons, impractical.
To address this limitation and to reduce the computational complexity of GED, several strategies for relaxing the problem have been proposed. These techniques range from binary linear programming (Justice and III, 2006), over fast suboptimal algorithms preserving classification accuracy (Neuhaus et al., 2006b; Riesen and Bunke, 2009a), to quadratic assignment models (Bougleux et al., 2017) and local search algorithms (Boria et al., 2020). More recently, diverse approaches to learn and approximate GED using Graph Neural Networks (GNNs) have been proposed (Bai et al., 2019a; Ranjan et al., 2022; Zhuo and Tan, 2022). A very recent approach (Jain et al., 2024), also based on GNNs, differs from the others in explicitly formulating the GED as a matrix optimization problem. That is, this method approximates the GED by learning the permutation that minimizes the assignment cost between the substructures of two graphs.
@@
Consequently, adapting GED to reflect meaningful, task-specific edit costs is non-trivial, since most available methods are supervised and rely on pre-computed distances. To make GED reflect functional similarity, it is necessary to learn the underlying edit costs – quantities that are not known a priori but constitute the very target of learning. While unsupervised methods such as GraphCL (You et al., 2020) or UGraphEmb (Bai et al., 2019b) can learn structural representations through augmentation, they cannot explicitly capture the actual edit costs. The problem of learning edit costs has a long history: early work pursued probabilistic and statistical formulations (Neuhaus and Bunke, 2004; 2007), while subsequent approaches relied on optimization strategies that are not fully differentiable end-to-end (Cortés et al., 2019; Garcia-Hernandez et al., 2020; Rica et al., 2021). More recent methods such as EPIC (Heo et al., 2024) and GMSM (Pellizzoni et al., 2024) introduce differentiable pipelines and attempt to learn edit costs. EPIC employs them primarily for data augmentation, while GMSM restricts node-specific costs to the final matching step, thereby missing multi-scale granularity.
@@ 2.2 Learning the Matching
In our model the matching between two graphs $G$ and $G^{\prime}$, containing $n$ and $m$ nodes respectively, is learned by constructing a series of matrices of size $n+m]\times[m+n]$, as illustrated in Figure 1 (a). A formal definition of these matrices is provided in Appendix B.
@@
The sum $\sum_{k=0}^{K}\mathbf{M}_{k}+\mathbf{D}+\mathbf{B}$ aligns with the LSAP-based reformulation of GED (Riesen and Bunke, 2009a). In our formulation, however, we do not constrain insertions and deletions to diagonal entries; instead, a masking mechanism hides off-diagonal elements from the network to prevent them from affecting learning (see $\mathbf{Mask}$ in Figure 1 (a)).
@@ 2.4 The Training of GEDAN
In the unsupervised setting, optimization follows Eq. 7 (see Appendix C.3 for the complete definition). If the weights of the functions $f_{\theta}^{(k)}$ are frozen and treated as fixed costs, GEDAN can be directly used to approximate GED, avoiding updates to the learned costs and preventing potential shortcut solutions. For consistency, the insertion and deletion costs ($a^{\ominus},a^{\oplus},b^{\ominus},b^{\oplus}$) are also kept fixed.
In the supervised setting, where target labels are provided, the model leverages both the edit costs produced by the functions $f_{\theta}^{(k)}$ and the scalar parameters $a^{\ominus},a^{\oplus},b^{\ominus},b^{\oplus}$ to align GED with the target labels. This gives GEDAN great flexibility. On the one hand, it can be used to minimize GED in a standard fashion using an appropriate loss (we use Smooth L1 loss (Girshick, 2015)). On the other hand, its most interesting application arises when topological dissimilarity does not coincide with functional dissimilarity. In this case, given the target labels, edit costs are optimized through a multi-task learning strategy (Song et al., 2016) that employs multiple loss functions. Multi-loss training in metric learning has been extensively studied (Wang et al., 2019a; Schroff et al., 2015) and is known to improve generalization and the robustness of learned representations. In our model, two complementary loss functions are employed.
@@
The second loss function is introduced to supervise the downstream task (Sobal et al., 2024), which may involve either regression or classification. This loss is computed via a neural network module that processes the predicted GEDs between the query graph and a set of key graphs, and uses this information to predict the label of the query graph. The complete definition of losses can be found in Appendix E.2.
@@ 3 Experimental Evaluation
For the first experiment, we use three datasets from the TUDataset repository (Morris et al., 2020) that contain molecule graphs. The chosen datasets consist of intentionally small graphs, motivated by the need to ensure the feasibility of the supervised training required by most models. For the second and third evaluations, we employ two benchmark datasets (Wu et al., 2017; Dwivedi et al., 2023b), namely FreeSolv (Mobley and Guthrie, 2014) and BBBP (Martins et al., 2012), which describe hydration free energy and blood–brain barrier permeability, respectively.
@@ 3.2 Optimization of Edit Costs
We use the predicted GED to embed the graphs in a vector space using dissimilarity-based embeddings (Riesen and Bunke, 2009b). The prototype selection (actually required for graph embedding) is performed uniformly to ensure a fair comparison. Moreover, to ensure the robustness of our results and to avoid possible bias due to an unrepresentative selection of prototypes, we repeat the process three times (for each fold during cross-validation).
@@ 3.3 Analysis of Edit Costs
In Figure 3 (c), we examine a molecule containing functional groups associated with enhanced water solubility. In particular, the sulfonyl (–SO2–) and nitro (–NO2) groups, which have been reported to influence solubility (Goldberg and Parker, 1985; Morrison and Boyd, 1992), are correctly highlighted by all models, albeit with different levels of resolution and local precision.
@@ 4 Conclusion
In this work, we propose a novel graph-based architecture that learns edit costs end-to-end, directly aligning GED with downstream tasks and overcoming the limitations of fixed-cost approaches. Although unsupervised GED approximation can improve graph matching in principle, achieving meaningful alignment with task-specific functional dissimilarity requires learning the edit costs. Our method enables fine-grained cost optimization at the node and subgraph level, producing embeddings that capture task-relevant differences and highlighting functionally important regions within the graphs. This fine-grained alignment not only improves predictive performance but also supports interpretable analysis by revealing which graph regions contribute most to the task.
Future work will focus on scaling the approach to larger graphs, extending cost learning to other graph types, and further investigating the theoretical properties of learned costs, aiming to provide bounds or monotonicity guarantees.
@@ Appendix C Model Details
Training GEDAN requires substantial computational resources. We employ a NVIDIA A100 GPUs for GED approximation, while experiments on learning edit costs are conducted on a cluster with two NVIDIA H100 GPUs to optimize training time and cost. The implementation is based on PyTorch (v2.2.0) together with the PyTorch Geometric library (v2.4.0), with hardware acceleration provided by CUDA (v12.1). The source code is publicly available at https://github.com/gedan-iclr2026/GEDAN.
@@ C.4 Gumbel-Sinkhorn
To obtain a good $\mathbf{P}$, we modify the version of the original Gumbel-Sinkhorn network (Mena et al., 2018; Shah et al., 2024), using the Truncated Gumbel (Neamah and Qasim, 2021) to reduce the gradient variance and add stability. We then sample with the Gumbel-Max Trick (Huijben et al., 2023) while maintaining differentiability. Finally we stabilize the normalization of Sinkhorn (Sinkhorn, 1964) with an early stop, improving efficiency and limiting gradient vanishing (LeCun et al., 1998).
@@ C.5 Influence of Depth and Inference Dynamics
This improvement is achieved in an unsupervised way, without direct supervision on the GED values, showing that $\text{GEDAN}(\lambda=1)$ can internally structure representations that aligns with true graph distances. Further examples are shown in Figures 131415 and 16.
@@ D.1 The Setting Up of the Configurations
    The first configuration serves as a baseline, where all operations – node and edge insertions, deletions, and substitutions – have uniform costs set to 1, as commonly assumed in several models.
@@
For completeness, we report the statistics of GED values for each configuration and dataset considered. Recent work indicates that GED benchmarks may have some bias related to isomorphism, or that this bias may be easily introduced during the training phase (Roy et al., 2025).
In our case, since the main goal is to compare our model with those in the literature, we explicitly address the risk of structural leakage across datasets. To mitigate this issue, we verify that the GED is equal to zero only for pairs of identical graphs. Concretely, we inspect the full GED matrix and ensure that zero values appear exclusively along the diagonal. We then exclude from the analysis any graph that exhibits a zero GED with a different graph (i.e., extra-diagonal zeros). Figure 9 shows the resulting GED matrix for our three datasets.
(a) The GED value matrices in the AIDS dataset
(b) The GED value matrices in the MUTAG dataset
(c) The GED value matrices in the PTC MR dataset
Figure 9: The matrices of GED values for configurations 1, 2, and 4, showing where the GED values are equal to zero.
@@ D.2 Results For Each Configuration
A comparative analysis across different configurations confirms the absence of a universally superior model, in agreement with the no-free-lunch theorem (Wolpert and Macready, 1997). The best performance is highlighted in boldface, while the second-best is underlined. For EUGENE (Bommakanti et al., 2024), we report the standard derivation, as we applied folding to accelerate the approximation using fewer graphs. In contrast, for the approximation proposed by Fischer et al. (2017), we approximated the entire dataset directly, since this method is efficient and, like EUGENE, does not require any training.
@@ Appendix E Details of Edit Cost Optimization
In this section, we provide additional details on the learning of edit costs. We illustrate the flexibility of the learnable cost functions employed by GEDAN, describe the datasets and loss functions used, and report further details on the results that are not covered in the main paper.
E.1 Details on Edit Cost Datasets
@@
FreeSolv (Mobley and Guthrie, 2014) BBBP (Martins et al., 2012) N. graphs Avg. nodes Avg. edges N. graphs Avg. nodes Avg. edges 642 8.72 16.78 1734 20.84 44.88
FreeSolv (Mobley and Guthrie, 2014) is a small dataset that collects hydration free energies (HFE), both experimental and calculated, for small molecules in water. It is commonly used for the prediction of molecular properties in the field of computational chemistry. The choice of this dataset is due to its small size, especially in application contexts one does not have a lot of data, so we verified that the model would work in this context.
@@
We introduce a new dataset, the GPU Network Cost Dataset (GNCD), specifically designed for our experiments. The dataset consists of synthetic graphs in which each node represents a GPU compute node, associated with its purchase cost, and each edge represents a cable required to connect two GPUs. The total value assigned to a graph corresponds to the overall cost of building the network, capturing the intrinsic cost asymmetry of the domain: GPU nodes are significantly more expensive than connection cables.
In our experimental setup, GNCD serves as a strongly asymmetric dataset for studying graph transformation costs. We define the transformation cost between two graphs as the cost of the components that appear in one graph but not in the other, with one important convention: if $G_{1}\subseteq G_{2}$, then the cost of transforming $G_{2}$ into $G_{1}$ is zero, since all components needed to obtain $G_{1}$ are already contained in $G_{2}$. This convention introduces a marked asymmetry not only between the costs associated with nodes and edges, but also between addition and deletion operations: deletion is considered free, whereas addition incurs a significant cost. We employ GNCD in the experiments reported in Section E.3.
The dataset contains 200 unique graphs, with an average of 13.2 nodes and 55.8 edges per graph. We include five different types of GPU computation nodes, priced at 2,400, 5,000, 6,000, 15,000, and 25,000 cost units, while edges (cables) cost only 25 units.
E.2 Loss and Training Strategy
@@
E.3 Learning Edit Costs Across Levels of Expressiveness
As defined by Formula 8 in GEDAN, cost learning is highly flexible, as it depends on the choice and implementation of the functions $f_{\theta}^{(k)}$. In this section, we analyze how GEDAN learns edit costs using datasets for which ground-truth edit costs are known. This setting simplifies the analysis compared to molecular datasets – where no such ground truth exists – and allows us to clearly illustrate how a framework like GEDAN can adapt to different domains. We examine three cost-function configurations:
  1. 1.
    Generic costs, independent of node type.
  2. 2.
    Node-specific costs combined with a linear function $f_{\theta}^{(k)}$, yielding a scalar-weighted sum that trades off expressiveness for linearity.
  3. 3.
    The original non-linear formulation combined with the GAM.
We conduct our analysis on three datasets. For MUTAG, we consider configurations 4 and 5 both asymmetric (see Table 6). In configuration 4, node and edge insertions cost twice as much as deletions, whereas in configuration 5 the opposite holds. The GNCD dataset, instead, is a strongly asymmetric collection of GPU-computation graphs, where deletions are free and insertions incur a cost. Node-specific GPU prices range from 2.4k to 25k units, and edges cost 25 units, providing a setting with heterogeneous node-dependent edit costs.
We begin with the generic-cost setting, where the functions $f_{\theta}^{(k)}$ are disabled. In this case, GEDAN can only learn the global edit costs $a^{\oplus},a^{\ominus},b^{\oplus},b^{\ominus}$, without distinguishing node types. In our experiment on MUTAG configurations 4 and 5, we prevent GEDAN from using the matrices $\mathbf{C}_{k}$, constraining it to learn only the generic edit parameters, which are randomly initialized to identical values across both configurations. Minimizing the training loss therefore forces GEDAN to adjust these generic costs toward their respective theoretical ground-truth values.
Figure 17: Trajectories of the edit costs $a^{\oplus},a^{\ominus},b^{\oplus},b^{\ominus}$ during optimization on MUTAG.
As shown in Figure 17, tracking the evolution of the costs during training reveals that GEDAN converges to an insertion-deletion ratio of 2:1 for configuration 4 and 1:2 for configuration 5. This happens despite the identical random initialization of $a^{\oplus},a^{\ominus},b^{\oplus},b^{\ominus}$, and matches the expected theoretical ratios: in configuration 4, node and edge insertions cost twice as much as deletions, whereas in configuration 5 deletions cost twice as much as insertions.
Figure 18: Cost trajectories during optimization on GNCD. (a) generic costs only; (b) linear fine-grained costs; (c) non-linear fine-grained costs.
In the second experiment, we compare the three cost-function configurations on the GNCD dataset. In Figure 18, we show the trajectories of both the generic and the fine-grained substitution costs, focusing specifically on node-insertion costs at the first level, i.e., $f_{\theta}^{0}(\epsilon,n_{i})$ where $n_{i}$ is one of the five node types in GNCD. Since ground-truth values are available, we can directly inspect whether the learned functions maintain the correct cost proportions at a finer granularity.
Table 16: Performance of the three GEDAN cost-function variants on MUTAG (conf. 4) and GNCD, showing how expressiveness affects accuracy.
GEDAN Edit Cost Datasets Generic Costs Scalar-Weighted Sum Non-linear (GAM) MUTAG (Conf. 4) $\operatorname{RMSE}~(\downarrow)$ 5.23 $\pm$ 0.99 4.93 $\pm$ 0.69 4.82 $\pm$ 0.58 $\tau~(\uparrow)$ 0.484 $\pm$ 0.15 0.514 $\pm$ 0.11 0.599 $\pm$ 0.10 $\rho~(\uparrow)$ 0.727 $\pm$ 0.09 0.742 $\pm$ 0.08 0.825 $\pm$ 0.06 GNCD (Appendix E.1) $\operatorname{NRMSE~\%}~(\downarrow)$ 11.52 $\pm$ 1.25 6.32 $\pm$ 0.67 4.96 $\pm$ 0.41 $\tau~(\uparrow)$ 0.379 $\pm$ 0.08 0.683 $\pm$ 0.04 0.762 $\pm$ 0.02 $\rho~(\uparrow)$ 0.521 $\pm$ 0.07 0.868 $\pm$ 0.03 0.944 $\pm$ 0.02
As shown in Table 16, the choice of cost function directly affects model performance. On MUTAG – where the task-specific functional distance $d_{f}$ is asymmetric but structurally simple – the three variants perform comparably, although the non-linear model consistently yields the best results.
In contrast, GNCD exhibits strong node-type–specific asymmetries, making generic-cost models underfit due to insufficient expressive capacity. Interestingly, all three models converge the generic node deletion cost toward zero during training, which is desirable since deletions are defined as free in GNCD. The main differences arise in the modeling of substitution costs: despite the well-defined cost structure, the linear model fails to represent these values accurately, whereas the non-linear GAM-based variant provides substantially more precise estimates, which is reflected in the better final metrics.
Figure 19: Predicted vs. ground-truth edit costs on GNCD for the three $f_{\theta}^{(k)}$ (generic, linear scalar-weighted, and non-linear GAM), illustrating how well the learned costs preserve the ordering of the true functional distances $d_{f}$.
E.4 Edit Cost Optimization Effect